378 𝑙𝑝𝑠 of air is isothermally compressed at 27℃ and 1379 π‘˜π‘ƒπ‘Žπ‘Ž; 𝑝2 = 4136 π‘˜π‘ƒπ‘Žπ‘Ž. For both nonflow and steady-flow (𝑉1 = 23 π‘š 𝑠, 𝑉2 = 46 π‘š 𝑠, βˆ†π‘ƒ = 0) processes, compute (a) ∫ 𝑝𝑑𝑉 π‘Žπ‘›π‘‘ βˆ’ ∫ 𝑉𝑑𝑝 (b) βˆ†π‘ˆ, βˆ†π», π‘Žπ‘›π‘‘ βˆ†π‘† (c) π‘Š π‘Žπ‘›π‘‘ 𝑄



1
Expert's answer
2021-10-31T18:12:24-0400

a)

"\int pdV =1379\cdot 10^3\displaystyle{\int^{46}_{23}}dV=1379\cdot 10^3\cdot 23=31717\cdot 10^3" J


"\int Vdp =378\displaystyle{\int^{4136}_{1379}}dp=378\cdot 2757=1042146" J


b)

"\Delta U=Q-W=0"


for non-flow:

"\Delta H=V\Delta p=1042146" J


for steady-flow:

"\Delta H=p\Delta V=31717\cdot 10^3" J


"\Delta S=Q/T"

"T=27-273=-246" K


for non-flow:

"Q=\int pdV=31717\cdot 10^3" J

"\Delta S=-31717\cdot 10^3/246=-129" J


for steady-flow:

"Q=\int Vdp =1042146" J

"\Delta S=-1042146/246=-4236" J


c)

for non-flow:

"Q=W=\int pdV=31717\cdot 10^3" J


for steady-flow:

"Q=W=\int Vdp=1042146" J